# Inverse Concave-Utility Reinforcement Learning is Inverse Game Theory

### v2

@@ Abstract
We consider inverse reinforcement learning problems with concave utilities. Concave Utility Reinforcement Learning (CURL) is a generalisation of the standard RL objective, which employs a concave function of the state occupancy measure, rather than a linear function. CURL has garnered recent attention for its ability to represent instances of many important applications including the standard RL such as imitation learning, pure exploration, constrained MDPs, offline RL, human-regularized RL, and others. Inverse reinforcement learning is a powerful paradigm that focuses on recovering an unknown reward function that can rationalize the observed behaviour of an agent. There has been recent theoretical advances in inverse RL where the problem is formulated as identifying the set of feasible reward functions. However, inverse RL for CURL problems has not been considered previously. In this paper we show that most of the standard IRL results do not apply to CURL in general, since CURL invalidates the classical Bellman equations. This calls for a new theoretical framework for the inverse CURL problem. Using a recent equivalence result between CURL and Mean-field Games, we propose a new definition for the feasible rewards for I-CURL by proving that this problem is equivalent to an inverse game theory problem in a subclass of mean-field games. We present initial query and sample complexity results for the I-CURL problem under assumptions such as Lipschitz-continuity. Finally, we outline future directions and applications in human–AI collaboration enabled by our results.
@@ 1 Introduction
Imagine that we are trying to infer the reward function of a human who has information-processing bounds, which is a form of bounded rationality due to cognitive constraints [26, 27, 28, 29]. This setting is often modelled as information-limited decision-making that is equivalent to the cautious / human-regularised RL formulation in table LABEL:tbl:table [30, 31, 32]. Here, $d^{\rho}$ represents a prior, low mental-effort policy for the human, and they pay a cost for deviating from it proportional to the KL divergence. The figure 1 presents an illustrative example, where a captain must navigate to a port  [^1] The shark, ship, and port icons created by Gilberto, hartadidesign, and Darwin Mulya from the Noun Project. . The captain’s true reward function is such that each time-step has a small cost, crossing into dangerous waters (represented as sharks) have a high cost, and the rewards at the ports induce a preference ordering so that the port A is preferred the most, and C the least. The environment is not stochastic, thus actions are always executed correctly. Here, the obviously optimal policy for the captain is to deterministically follow the golden path. However, being under a heavy cognitive burden, the captain knows they are prone to making errors. Thus instead, they follow a more forgiving path to port B, which leaves room for error. In similar episodes, they do make mistakes, and we end up observing a dataset of demonstrations within the green-shaded region. This is not the same as having a stochastic environment, since another captain might successfully execute the golden path in the same environment, and we already know the dynamics are deterministic. Here, even in the limit of infinite demonstrations, we cannot infer the true reward function with standard IRL. We prove in lemma 7 that indeed the true reward function of a CURL problem may never be inferred via standard IRL even in the case of known expert policy.
@@
In section 3, we prove how the feasible reward set definition of standard IRL fails in inverse CURL (I-CURL). Then using the equivalence to MFGs, we propose a new set, and prove that it is sufficient and necessary, containing all rewards that rationalise the optimal CURL policy. We show an equivalent characterization of this set where computing a feasible reward function is solving a constant-sum game itself. The rest of this section analyses two different cases: (1) Given the expert policy $\pi^{E}$, under what assumptions can we compute a feasible reward function efficiently, and (2) Given an empirical estimate of the expert policy $\hat{\pi}^{E}$ from a dataset, what is the relationship between the sets of feasible rewards for $\hat{\pi}^{E}$ and $\pi^{E}$. Finally, we derive a sample complexity result for the case (2). The section 3.3 examines some of the assumptions we made for our theoretical results, and shows that in practice they hold for many CURL problems. Since the inverse CURL problem has not been addressed before, our work fills an important theoretical gap, demonstrating the limitations of standard IRL methods for I-CURL. Our formulation enables the inference of reward functions from bounded rational behaviour such as information-boundedness and risk-aversion.
@@ 3.2 Empirical I-CURL
Assumption 13
The CURL-MFG reward function $R(\cdot,\cdot,d)$ is $L$-Lipschitz continuous so that $||R(\cdot,\cdot,d^{\prime})-R(\cdot,\cdot,d)||_{\infty}\leq L||d^{\prime}-d||_{\infty}.$
Since the true CURL-MFG reward function is assumed to be Lipschitz-continuous, here we will restrict $\mathcal{R}_{\mathbb{B}}$ and $\mathcal{R}_{\hat{\mathbb{B}}}$ to $L$-Lipschitz functions. An important relaxation of the Nash equilibrium is the $\epsilon$-Nash equilibrium, defined as any policy $\pi$ that achieves a return that is $\epsilon$ close to the return of a Nash policy, so that its exploitability satisfies the inequality $\phi(\pi;\theta)\leq\epsilon.$ The following theorem establishes the symmetric relationship between $\mathcal{R}_{\mathbb{B}}$ and $\mathcal{R}_{\hat{\mathbb{B}}}$ in terms of $\epsilon$-equilibria.
Theorem 14
Under the assumption 13, let $||d^{E}-\hat{d}^{E}||_{\infty}\leq\epsilon.$ Then for any $R\in\mathcal{R}_{\mathbb{B}},$ $\hat{\pi}^{E}$ is a $2L\epsilon$-equilibrium. Similarly, for any $\hat{R}\in\mathcal{R}_{\hat{\mathbb{B}}},$ the $\pi^{E}$ is a $2L\epsilon$-equilibrium.
Proof. Let $R\in\mathcal{R}_{\mathbb{B}}$ and thus $\phi(\pi^{E};R)=0.$ Then $|\phi(\hat{\pi}^{E};R)-\phi(\pi^{E};R)|$ can be written as $|\max_{d\in K_{\gamma}}\langle d,R(\cdot,\cdot,\hat{d}^{E})\rangle-max_{d\in K_{\gamma}}\langle d,R(\cdot,\cdot,d^{E})\rangle-\langle\hat{d}^{E},R(\cdot,\cdot,\hat{d}^{E})\rangle+\langle d^{E},R(\cdot,\cdot,d^{E})\rangle|.$ Under the assumption 13, we have that $|\langle d^{E},R(\cdot,\cdot,d^{E})\rangle-\langle\hat{d}^{E},R(\cdot,\cdot,\hat{d}^{E})\rangle|\leq L||\hat{d}^{E}-d^{E}||_{\infty}.$ Same clearly holds for the difference between max terms. If $||\hat{d}^{E}-d^{E}||_{\infty}\leq\epsilon,$ then we have the inequality $|\phi(\hat{\pi}^{E};R)-\phi(\pi^{E};R)|\leq 2L||\hat{d}^{E}-d^{E}||_{\infty}\leq 2L\epsilon.$ Since $\phi(\pi^{E};R)=0$ and $\phi(\pi;R)\geq 0$ for any $\pi$ by definition, we have $\phi(\hat{\pi}^{E};R)\leq 2L\epsilon.$ Therefore $\hat{\pi}^{E}$ is a $2L\epsilon$-equilibrium for reward $R.$ Similarly, take $\hat{R}\in\mathcal{R}_{\hat{\mathbb{B}}}$ such that $\phi(\hat{\pi}^{E};\hat{R})=0.$ We can do a similar derivation with $\phi(\pi^{E};\hat{R})-\phi(\hat{\pi}^{E};\hat{R})$ and end up with $\phi(\pi^{E};\hat{R})\leq 2L\epsilon.$ Therefore $\pi^{E}$ is an $2L\epsilon$-equilibrium for reward $\hat{R}.$
We will use this symmetry to show how the similarity between $\mathcal{R}_{\mathbb{B}}$ and $\mathcal{R}_{\hat{\mathbb{B}}}$ can be measured. Namely, we propose that two CURL-MFG reward functions are similar if the equilibrium of one is an $\epsilon$-equilibrium of the other. Such reward functions are sometimes called $\epsilon$-inverse equilibria [40]. Important to note that the inequalities above hold for any pair of rewards $R\in\mathcal{R}_{\mathbb{B}},\hat{R}\in\mathcal{R}_{\hat{\mathbb{B}}}.$ Then we will say that the dissimilarity between the two sets is $D(\mathcal{R}_{\mathbb{B}},\mathcal{R}_{\hat{\mathbb{B}}})=2L\epsilon.$ It is clear that if the gap $||d^{E}-\hat{d}^{E}||_{\infty}$ gets smaller, the two sets become more similar since $\epsilon$ gets smaller. This means within the class of $L$-Lipschitz reward functions, we can control the I-CURL error by the estimation error. The following lemma presents a sample complexity bound to achieve $\leq\epsilon$ error with high probability.
Lemma 15 (Sample Complexity of Estimating $d^{E}$)
Let $n$ be the number of samples from the equilibrium $(\pi^{E},d^{E})$ such that $(s_{i},a_{i})\sim d^{E}(s,a)$ for $i\in[n]$. Then for every $\epsilon>0,$ the estimation error can be bounded as $Pr(||d^{E}-\hat{d}^{E}||_{\infty}>\epsilon)\leq 2e^{-2n\epsilon^{2}}.$
Proof. The proof is a direct application of a so-called "folklore fact" about learning discrete distributions (see Theorem 10 of [41]), which itself is due to the classical result of Dvoretzky et al. [42] and the fact that $||d^{E}-\hat{d}^{E}||_{\infty}\leq 2d_{k}(d^{E},\hat{d}^{E})$ where $d_{k}$ is the Kolmogorov distance.
The following corollary is a structural result that follows directly from definitions, and tells us how the dynamics $T$ and the initial state distribution $\mu_{0}$ affect the difficulty of the I-CURL problem. For example, if the starting state is fixed in a chain-like dynamics with low stochasticity, the diameter $D$ can be really low. Then any policy $\pi$ will induce very similar occupancy measures. This means we might want a more fine-grained error control with an $\epsilon$ that is relative to $LD$ in the future.
Corollary 16 (The Effect of $K_{\gamma}$’s Diameter)
Under the assumption 13, let $D=\max_{d^{\prime},d\in\mathcal{K}_{\gamma}}||d^{\prime}-d||_{\infty}$ denote the diameter of $\mathcal{K}_{\gamma}.$ Then $\phi(\pi;R)\leq 2LD$ for any $\pi.$
Proof. Follows from theorem 14 by replacing $\epsilon$ with $D.$
3.3 Discussion on the Theoretical Assumptions and Practical Applications
The assumption 13 is in general a reasonable assumption for a reward function, since even for continuous state-action spaces, bounded rewards is a common assumption in RL. This means we can always choose $L=R_{\text{max}}D$ to be the maximum possible reward times the diameter of the $K_{\gamma}.$ However, in this particular case, the CURL-MFG reward is defined as $\nabla F$, and the gradient of a concave function need not be finite everywhere. Below, we will show that in practice this is not an issue for many CURL problems.
Consider the setting where $F(d^{\pi})=(1-\lambda)\langle d^{\pi},R\rangle-\lambda KL(d^{\pi}||d^{\rho})$ with a fixed $d^{\rho}$. Depending on $\lambda$, this formulation can represent various CURL problems such as divergence minimization, state-marginal matching, offline RL, human-regularized RL, and cautious RL (see table LABEL:tbl:table). Here the induced CURL-MFG’s reward function is $R(\cdot,\cdot,d^{\pi})=(1-\lambda)\langle d^{\pi},R\rangle-\lambda\nabla[KL(d^{\pi}||d^{\rho})].$ Ignoring the first term which is linear and bounded, the assumption 13 holds if and only if $|\nabla KL(d^{\prime}||d^{\rho})-\nabla KL(d||d^{\rho})|\leq L^{\prime}||d-d^{\prime}||$ for some $L^{\prime}$ as well. Unfortunately, KL divergence does not satisfy this bound, since it is not Lipschitz smooth. Therefore, it appears to be that any CURL problem with a KL-divergence term in the objective $F$ fails to satisfy our assumptions.
The reason why KL-divergence is not Lipschitz smooth is due to the fact that its gradient can be infinite. Fortunately, in many practical CURL problems this is not the case, which means our assumptions are valid for them. Consider the case $R(\cdot,\cdot,d^{\pi})=\langle d^{\pi},R\rangle+\nabla[-\frac{1}{\beta}KL(d^{\pi}||d^{\rho})]$ where now $d^{\rho}$ and $d^{\pi}$ are the state occupancy measures. As long as $d^{\rho}$ and $d^{\pi}$ are absolutely continuous with respect to each other, the gradient of the KL divergence is guaranteed to be finite. One corollary that follows from this is the fact that ergodicity (i.e. every state has a non-zero occupancy for any policy) guarantees assumption 13 in these type of problems. In the context of offline RL, the $d^{\rho}$ can be a state-action occupancy measure estimated from a dataset. A common assumption in offline RL is the uniformly bounded concentrability coefficient, which means $\frac{d^{\pi}(s,a)}{d^{\rho}(s,a)}\leq C$ for some finite $C$ everywhere [43]. This assumption also ensures Lipschitz-smoothness of the KL term in finite state-action spaces. Similar arguments can be made for instance for constrained MDPs through smooth barrier functions such as Zhang et al. [44].
@@ 4 Related Works
The goal of population-level approaches is to infer the population-level reward of the form $R(d^{\pi},\pi)$ from a dataset consisting of occupancy measure and policy pairs $(d_{0},\pi_{0}),...,(d_{T},\pi_{T})$ instead of state-action tuples [54]. Our approach is based on individual-level inverse game theory for mean-field games, which has been explored empirically for general mean-field games [39, 55]. We focus particularly on leveraging the structure of MFGs induced by CURL problems, and derive finite-sample results. The problem of rationalising observed multi-agent behaviour by inferring rewards has been referred to as inverse multi-agent RL [56], inverse game theory [37], and inverse equilibrium problems [38] in different contexts. Recent formulation by Goktas et al. [40] represented the IGT problem itself as a zero-sum game whose solution set is referred to as the inverse Nash equilibria.
@@ 6 Conclusion
Standard inverse reinforcement learning is not directly applicable to inferring reward functions of agents that have optimized their policies for solving a concave utility reinforcement learning problem. Even in the idealized case of having access to the expert policy $\pi^{E}$, we have shown that the set of feasible reward functions as defined by the standard IRL literature can be empty. We have resolved this problem by proving a game-theoretic characterization of the feasible reward set for inverse CURL, leveraging a result that shows every CURL problem is equivalent to a mean-field game. Our theorem 12 shows that the exact computation under known $\pi^{E}$ can be tackled efficiently with the AIMP method of concurrent work under a convexity-concavity assumption [40]. In theorem 14, we have proven that for the class of L-Lipschitz reward functions, a better estimation of $d^{E}$ implies a feasible set more similar to $\mathcal{R}_{\mathbb{B}}$ in terms of the induced equilibria. This allows us to control the error of I-CURL via the estimation error, providing us with the sample complexity result in lemma 15. To our knowledge, our work presents the first formulation and theoretical analysis of the inverse CURL problem.

### v3

@@ Abstract
We consider inverse reinforcement learning problems with concave utilities. Concave Utility Reinforcement Learning (CURL) is a generalisation of the standard RL objective, which employs a concave function of the state occupancy measure, rather than a linear function. CURL has garnered recent attention for its ability to represent instances of many important applications including the standard RL such as imitation learning, pure exploration, constrained MDPs, offline RL, human-regularized RL, and others. Inverse reinforcement learning is a powerful paradigm that focuses on recovering an unknown reward function that can rationalize the observed behaviour of an agent. There has been recent theoretical advances in inverse RL where the problem is formulated as identifying the set of feasible reward functions. However, inverse RL for CURL problems has not been considered previously. In this paper we show that most of the standard IRL results do not apply to CURL in general, since CURL invalidates the classical Bellman equations. This calls for a new theoretical framework for the inverse CURL problem. Using a recent equivalence result between CURL and Mean-field Games, we propose a new definition for the feasible rewards for I-CURL by proving that this problem is equivalent to an inverse game theory problem in a subclass of mean-field games. We outline future directions and applications in human–AI collaboration enabled by our results.
@@ 1 Introduction
Imagine that we are trying to infer the reward function of a human who has information-processing bounds, which is a form of bounded rationality due to cognitive constraints [26, 27, 28, 29]. This setting is often modelled as information-limited decision-making that is equivalent to the cautious / human-regularised RL formulation in table LABEL:tbl:table [30, 31, 32]. Here, $d^{\rho}$ represents a prior, low mental-effort policy for the human, and they pay a cost for deviating from it proportional to the KL divergence. The figure 1 presents an illustrative example, where a captain must navigate to a port  [^1] The shark, ship, and port icons created by Gilberto, hartadidesign, and Darwin Mulya from the Noun Project. . The captain’s true reward function is such that each time-step has a small cost, crossing into dangerous waters (represented as sharks) have a high cost, and the rewards at the ports induce a preference ordering so that the port A is preferred the most, and C the least. The environment is not stochastic, thus actions are always executed correctly. Here, the obviously optimal policy for the captain is to deterministically follow the golden path. However, being under a heavy cognitive burden, the captain knows they are prone to making errors. Thus instead, they follow a more forgiving path to port B, which leaves room for error. In similar episodes, they do make mistakes, and we end up observing a dataset of demonstrations within the green-shaded region. This is not the same as having a stochastic environment, since another captain might successfully execute the golden path in the same environment, and we already know the dynamics are deterministic. Here, even in the limit of infinite demonstrations, we cannot infer the true reward function with standard IRL.
@@
In section 3, we prove how the feasible reward set definition of standard IRL fails in inverse CURL (I-CURL). Then using the equivalence to MFGs, we propose a new set, and prove that it is sufficient and necessary, containing all rewards that rationalise the optimal CURL policy. We show an equivalent characterization of this set where computing a feasible reward function is solving a constant-sum game itself. Since the inverse CURL problem has not been addressed before, our work fills an important theoretical gap, demonstrating the limitations of standard IRL methods for I-CURL. Our formulation enables the inference of reward functions from bounded rational behaviour such as information-boundedness and risk-aversion.
@@ 4 Related Works
The goal of population-level approaches is to infer the population-level reward of the form $R(d^{\pi},\pi)$ from a dataset consisting of occupancy measure and policy pairs $(d_{0},\pi_{0}),...,(d_{T},\pi_{T})$ instead of state-action tuples [54]. Our approach is based on individual-level inverse game theory for mean-field games, which has been explored empirically for general mean-field games [39, 55]. The problem of rationalising observed multi-agent behaviour by inferring rewards has been referred to as inverse multi-agent RL [56], inverse game theory [37], and inverse equilibrium problems [38] in different contexts. Recent formulation by Goktas et al. [40] represented the IGT problem itself as a zero-sum game whose solution set is referred to as the inverse Nash equilibria.
@@ 6 Conclusion
Standard inverse reinforcement learning is not directly applicable to inferring reward functions of agents that have optimized their policies for solving a concave utility reinforcement learning problem. Even in the idealized case of having access to the expert policy $\pi^{E}$, we have shown that the set of feasible reward functions as defined by the standard IRL literature can exclude the true reward function. We have resolved this problem by proving a game-theoretic characterization of the feasible reward set for inverse CURL, leveraging a result that shows every CURL problem is equivalent to a mean-field game.